Need help solving this dynamics problem
HW 7 (Section 15.5-15.7 & 16.1-16.5) Fundamental Problem 16.12 9 of 10 End A of the link has a velocity of vA 6 m/s 2 m 45 Part A Determine the velocity of the peg at B at this instant. The peg is constrained to move along the slot. Express your answer to three significant figures and include the appropriate units. UB-1 Value Units Submit Request Answer
Answer #1

The velocity polygon can be drawn as shown below:

solving the velocity polygon using sine rule we get;

\frac{V_{A}}{sin 15}=\frac{V_{B}}{sin 30}=\frac{V_{B/A}}{sin 135}

\frac{6}{sin 15}=\frac{V_{B}}{sin 30}

6 sin30 sin15

V_{B}=11.59 m/s .....Ans.